# Deep Learning; Physics-Informed Neural Networks; Partial Differential Equations; Schrödinger equation

### v1

@@ Abstract
Quantum many-body systems are of great interest for many research areas, including physics, biology and chemistry. However, their simulation is extremely challenging, due to the exponential growth of the Hilbert space with the system size, making it exceedingly difficult to parameterize the wave functions of large systems by using exact methods. Neural networks and machine learning in general are a way to face this challenge. For instance, methods like Tensor networks and Neural Quantum States are being investigated as promising tools to obtain the wave function of a quantum mechanical system. In this tutorial, we focus on a particularly promising class of deep learning algorithms. We explain how to construct a Physics-Informed Neural Network (PINN) able to solve the Schrödinger equation for a given potential, by finding its eigenvalues and eigenfunctions. This technique is unsupervised, and utilizes a novel computational method in a manner that is barely explored. PINNs are a deep learning method that exploits Automatic Differentiation to solve Integro-Differential Equations in a mesh-free way. We show how to find both the ground and the excited states. The method discovers the states progressively by starting from the ground state. We explain how to introduce inductive biases in the loss to exploit further knowledge of the physical system. Such additional constraints allow for a faster and more accurate convergence. This technique can then be enhanced by a smart choice of collocation points in order to take advantage of the mesh-free nature of the PINN. The methods are made explicit by applying them to the infinite potential well and the particle in a ring, a challenging problem to be learned by an AI agent due to the presence of complex-valued eigenfunctions and degenerate states
@@ 2 Materials and Methods
Figure 1: Graphical representation of the approach followed in this paper. The circles represent variables, the rectangle represent operations. In purple there are the two neural networks, the main one that computes the eigenstates and the one that computes the energy. In red there are the outputs of the neural networks: the auxiliary outputs, the main outputs which is the eigenfunction and the eigenvalue $E$. In yellow there are the operations needed to calculate the losses. The losses are the integral loss, the boundary condition loss, the normalization loss, the inductive biases and the differential equation loss. SAE is the Sum of Absolute Errors, SSE is the Sum of Squared Errors (Also see Eq. (5)) In green there is the final loss. It is calculated by summing the partial losses weighted by empirically adjusted hyperparameters $w_{int}$, $w_{bc}$, $w_{norm}$, $w_{bias}$ and $w_{eq}$. In cyan there are the criteria and the optimization method. The grey arrows represent the steps where numerical integration is required.
@@ 2.1 Neural Networks and Deep Learning
This Section gives a broad qualitative overview of the basic tools needed for this tutorial, namely deep neural networks. The information here has been largely taken from Refs. Kruse et al. (2022) and Alzubaidi et al. (2021). (Artificial) neural networks are information processing systems, whose structure and operation principles are inspired by the nervous system and the brain of animals and humans. They consist of a large number of fairly simple units, the so-called neurons, which are working in parallel. These neurons communicate by sending information in the form of activation signals, along directed connections, to each other. Deep Learning algorithms face a series of challenges to train them effectively. For instance, one of the most common issues is the problem of overfitting. With a larger number of hidden neurons, a neural network may adapt not only to the regular dependence between outputs and inputs we try to capture, but also to the accidental specifics (and thus also the errors and deviations) of the training data set. This is usually called overfitting. Overfitting will usually lead to the effect that the error a neural network with many hidden neurons yields on the validation data [^1] A second dataset whose labels are known that is not employed in training but only to get a more accurate estimate of the error of the network will be greater than the error on the training data. This is due to the fact that the validation data set is likely distorted in a different fashion, since the errors and deviations are random. Furthermore, while DL methods are known to be extremely effective when immense amounts of data are available, this availability is not a given in all circumstances. For instance, the data might be costly to obtain. Or it might be tied to the observation of some phenomenon that cannot be replicated at will and is not common enough to allow the training of a DL model that relies on big data. PINNs aim to addresso both the problem of data availability and mitigate overfitting by making use of the underlying physics of the system.
@@ 2.2 Materials
All results in this paper are reproducible using a laptop with a NVIDIA GeForce RTX 3060 Laptop GPU.
@@ 2.3 Physics-Informed neural networks
and the normalization will be fulfilled if $\nu(\frac{L}{2})=1$; $\nu(-\frac{L}{2})=0$. Note that there is a trade-off here. On the one hand the discretization error can be potentially removed using an auxiliary output. On the other hand using it makes the network harder to train due to the additional outputs and losses. Therefore the choice to utilize an Auxiliary output or a finite difference method has to be made on a case-by-case basis.
@@ 2.5 Physical constraints and inductive biases
After defining the main differential equation loss in the previous Section, in this Section the other losses used to build a neural network capable of solving the equation are explained. In particular, here are defined the losses linked to the physical constraints of the system. We also add a kind of loss not essential to solve the equation but which greatly improves convergence, namely the inductive biases.
To define these losses one needs to understand which are the physical constraints of the system and if any symmetries can be exploited to obtain an inductive bias. By inductive bias we mean a set of conditions that leads the network to prioritize a set of solution over other, for instance prioritizing solutions that follow a given symmetry. Biases can be invaluable because they limit the search space, and thus reduce the number of epochs needed to obtain a solution.
@@
where $\Omega$ is the domain of the wave function, chosen depending on the dimensionality and nature of the system. For instance, for a one-dimensional infinite potential well in $[-\frac{L}{2},\frac{L}{2}]$, $\Omega$ can be restricted to $\Omega=[-\frac{L}{2},\frac{L}{2}]$. It is important to note that imposing the condition defined in (11) is the only thing that forces the wave function to be nonzero. Otherwise $\psi(x)=0$ for all $x$ would exactly satisfy all losses. This integral can be solved by utilizing an auxiliary output $\nu(x)$ with form (8) and by imposing $\nu(\frac{L}{2})=1$; $\nu(-\frac{L}{2})=0$. However, we must first train the network so that $\nu(x)$ has the desired form (8). Once again, this can be achieved by exploiting automatic differentiation, to compute $\gradient\nu(\vec{x})$ and thus use the following as a loss:
@@
obtaining the integral loss. We typically need to set some boundary conditions. In general, one can either use hard or soft boundary conditions. The former choice forces the network output to fulfill the boundary conditions exactly. This can be done by employing a change of variable or modifying in some way the output, such as by multiplying it by another function. The latter choice makes the network learn the BCs via an additional loss, the boundary condition loss. This loss is 00 when the BCs are fulfilled. For instance in the infinite potential well to impose soft Dirichlet Boundary conditions (10) we can add to the network’s loss $|\psi(-\frac{L}{2})|+|\psi(\frac{L}{2})|$. In this tutorial we always use soft boundary conditions. Further constraints can be required on a case-by-case basis, but normalization and boundary conditions are always essential. Moreover, we can also employ an additional kind of loss: the inductive biases. These are losses that are not required to obtain a solution, but can instead be added to allow the network to converge much faster. They are usually based on the symmetries of the potential: for instance, the infinite well is symmetric with respect to 00. This symmetry makes its eigenfunctions alternatively even and odd with respect to 00. A conditions that can be enforced in the PINN by the symmetry loss:
@@
where $s$ is $1$ for symmetric states and $-1$ for antysimmetric ones (the value of $s$ is switched each time a network is trained, starting from $s$=1).
@@ 2.6 Self-consistent computation of the eigenenergy
In this Section, we explain how to find the eigenvalue to pair with the eigenfunction obtained thanks to the techniques described in the previous Sections. Currently, the network can only find the eigenfunction of the potential when the energy is known. However, in general when solving the Schrödinger equation that will not be the case. Therefore a way to find it in parallel with the eigenfunction is needed. There are two ways to do this: the first is to set the energy $E$ itself as a trainable parameter, but this seems to result in fairly slow training. The second is to build a second neural network to train in parallel with the one that predicts the wave function. It will take as input a constant stream of "$1"$s of the same length as the input vector, and output a guess for the energy $E$. Since the output is constant, this second network will have no hidden layers and there will be no nonlinear activation function. The network will just perform a linear transformation from $1$ to the value of the energy, based on a trainable parameter.
@@ 2.7 Finding the desired state
A crucial step is to build the losses that allow the network to distinguish the desired eigenstate from the infinite spectrum of the Hamiltonian. This is explained in this Section. By training the PINN with the current losses, it will solve the Schrödinger equation for a random state anywhere in the spectrum. Usually, however, we are looking for some amount of eigenstate starting from the ground state and going in order of increasing energy. In particular it is especially important to build a network that finds the ground state first. This can be achieved by introducing an energy minimization condition in the loss:
@@
where $E_{PINN}$ is the predicted energy, $a$ is a hyperparameter. $E_{init}$ is either another hyperparameter when training the network to predict the ground state, or the energy of the previous state, for the excited states. The choice to increase $E_{init}$ as the quantum number increases has been made to avoid the loss exploding for higher energy states due to its exponential increase. Otherwise the network might fail to converge at all. We also have to take into account that this energy minimization loss keeps pushing the network to lower energies even when the correct solution is reached. This moves the loss’ global minima from the eigenstate to a lower energy state, causing the loss to converge to an incorrect state. To avoid this the loss’ weight is gradually decreased as the training continues. This reduction should be such that the loss becomes approximately 00 towards the end of the training. Thanks to this in the earlier epochs this loss will steer the network towards the desired lower energy state. On the other hand in the latter epochs the energy minimization loss will become negligible and allow the network to settle to the correct solution. The resulting neural network can now find the ground state of the wave function. We now want to obtain the excited states. To do so, a sequence of network is built, each of which predicts one of the Hamiltonian’s eigenstates and the corresponding eigenvalue, in order of increasing energy. To do so, an additional loss is employed, the orthogonality loss. This loss forces each network’s output to be orthogonal to all the ones before it:
@@ 2.8 Weights
In the previous Sections the losses to minimize have been defined. In the following Sections is explained how to build the actual computational framework that ensures the network is able to minimize those losses. In this Section, in particular, some pointers are given on how to tune one of the most important set of hyperparameters: the weights. Usually, when training PINNs it is better to set weights that ensure all the losses have the same scale. That is not the case for this problem: it is very easy for a network that has to predict eigenfunctions to get stuck in local minimas such as $\psi(\vec{x})=0$ $\forall\vec{x}$ or an already discovered eigenvector. This means that it is actually better to set the weights in such a way that the network first enforces the physical constraints of the system, such as normalization and orthogonality. Only once these are fulfilled the network should be trained to actually solve the equation. This means setting the losses for those physical constraints higher that the differential equations’. In general, this means than the integral loss should have the highest weight, then the normalization, needed to avoid the constant solution, and the orthogonality, needed to avoid repeating solutions. At this point the network should minimize the boundary conditions loss. The last loss to minimize is the differential equation loss. However, since solving the differential equation is the main objective of the network it is often worth it to gradually increase its loss’ weight. This allows the network to learn how to enforce the physical conditions in the first epochs but makes it focus more on the differential equation in the latter epochs. It is important to stress that weights are probably the most impactful hyperparameter to tune: a correct choice of weights will be the difference between quick convergence and a network getting stuck on a suboptimal solution.
@@ 2.9 Choice of training points
In order to train the PINN, the losses must be evaluated on a set of points, known as collocation points. These can be taken anywhere in the domain, and this Section contains the optimal way to choose them. When training a PINN such as this, each batch of training points should span the whole domain. Furthermore, since no dataset is being used, the choice of points is not constrained in any way: they can be taken anywhere in the domain. This can be exploited to basically work with an infinitely large dataset. To fully take advantage of this, first a batch size $N$ is decided. Then a mesh of $N$ points that spans the whole domain is built. Each of this points will be the centre of a normal distribution from which one of the points in the given batch will be sampled. This results in batches that are always slightly different but always span the whole domain, thus ensuring the maximum exploration and fastest possible convergence.
@@ 2.11 Network architecture
In this Section some suggestions are given on how to better build a feed-forward neural network to solve the Schrödinger in a physics-informed way. We focus on which metrics to use for the loss and which activation function should be employed in the network. PINNs tend to work better as relatively shallow but wide networks. For instance in Brevi et al. (2024) the main network was made up of $7$ hidden layers of $256$ neurons each. As for the activation function, an infinitely differentiable one such as $tanh$ is needed. Otherwise the network would be unable to compute higher order derivative via automatic differentiation. Choosing an activation function of this type instead of something like $ReLU$ might, however, cause vanishing gradient problems. If this happens, it might be worth it to look into adaptive activation functions Jagtap et al. (2019) as a solution. Lastly, it is not necessary to compute all the partial losses using the same metric. It usually makes more sense to utilize something like Sum of Absolute Errors for the losses enforcing the physical conditions of the system, such as the Boundary Conditions loss or the Normalization loss. This is because the rest of the training makes sense only if those losses are approximately 00 in all points. And this is enforced more strictly with SAE rather then something like MSE. On the other hand, Mean Squared Error is usually the best choice for the integral and differential equation losses. Utilizing Sum of Squared Error to compute these losses is also worth considering. This especially true when the wave function is expected to have a complex behaviour in a small section of the domain and a relatively simple one in the rest. This choice of metric will avoid underestimating the loss due to the averaging of the error for the points in the large "simple" area and the small "complex" area. As for the optimizer, Adam Kingma and Ba (2015) is usually rather effective.

### v2

@@ Abstract
Quantum many-body systems are of great interest for many research areas, including physics, biology and chemistry. However, their simulation is extremely challenging, due to the exponential growth of the Hilbert space with the system size, making it exceedingly difficult to parameterize the wave functions of large systems by using exact methods. Neural networks and machine learning in general are a way to face this challenge. For instance, methods like Tensor networks and Neural Quantum States are being investigated as promising tools to obtain the wave function of a quantum mechanical system. In this tutorial, we focus on a particularly promising class of deep learning algorithms. We explain how to construct a Physics-Informed Neural Network (PINN) able to solve the Schrödinger equation for a given potential, by finding its eigenvalues and eigenfunctions. This technique is unsupervised, and utilizes a novel computational method in a manner that is barely explored. PINNs are a deep learning method that exploits Automatic Differentiation to solve Integro-Differential Equations in a mesh-free way. We show how to find both the ground and the excited states. The method discovers the states progressively by starting from the ground state. We explain how to introduce inductive biases in the loss to exploit further knowledge of the physical system. Such additional constraints allow for a faster and more accurate convergence. This technique can then be enhanced by a smart choice of collocation points in order to take advantage of the mesh-free nature of the PINN. The methods are made explicit by applying them to the infinite potential well and the particle in a ring, a challenging problem to be learned by an Artificial Intelligence agent due to the presence of complex-valued eigenfunctions and degenerate states
@@ 2 Materials and Methods
Figure 1: Graphical representation of the approach followed in this paper. The circles represent variables, the rectangles represent operations. The variables are also represented by primary colors, while the methods are represented by the secondary colors obtained by combining the colors of their input and their output. In green there are the neural network inputs. In yellow there are the two neural networks, the main one that computes the eigenstates and the one that computes the energy. In blue there are the outputs of the neural networks: the auxiliary outputs, the main outputs which is the eigenfunction and the eigenvalue $E$. In purple there are the operations needed to calculate the losses. The losses are the integral loss, the boundary condition loss, the normalization loss, the inductive biases and the differential equation loss. SAE is the Sum of Absolute Errors, SSE is the Sum of Squared Errors. In red there is the final loss. It is calculated by summing the partial losses weighted by empirically adjusted hyperparameters $w_{int}$, $w_{bc}$, $w_{norm}$, $w_{bias}$ and $w_{eq}$. In cyan there are the criteria and the optimization method.
@@ 2.1 Neural Networks and Deep Learning
This Section gives a broad qualitative overview of the basic tools needed for this tutorial, namely deep neural networks. The information here has been largely taken from Refs. Kruse et al. (2022) and Alzubaidi et al. (2021). (Artificial) neural networks are information processing systems, whose structure and operation principles are inspired by the nervous system and the brain of animals and humans. They consist of a large number of fairly simple units, the so-called neurons, which are working in parallel. These neurons communicate by sending information in the form of activation signals, along directed connections, to each other. Deep Learning algorithms face a series of challenges to train them effectively. For instance, one of the most common issues is the problem of overfitting. With a larger number of hidden neurons, a neural network may adapt not only to the regular dependence between outputs and inputs we try to capture, but also to the accidental specifics (and thus also the errors and deviations) of the training data set. This is usually called overfitting. Overfitting will usually lead to the effect that the error a neural network with many hidden neurons yields on the validation data [^1] A second dataset whose labels are known that is not employed in training but only to get a more accurate estimate of the error of the network will be greater than the error on the training data. This is due to the fact that the validation data set is likely distorted in a different fashion, since the errors and deviations are random. Furthermore, while DL methods are known to be extremely effective when immense amounts of data are available, this availability is not a given in all circumstances. For instance, the data might be costly to obtain. Or it might be tied to the observation of some phenomenon that cannot be replicated at will and is not common enough to allow the training of a DL model that relies on big data. PINNs aim to address both the problem of data availability and mitigate overfitting by making use of the underlying physics of the system.
@@ 2.2 Materials
Due to the pedagogical nature of this article, we decided to not use examples that require expensive computational resources. Therefore, all the results in the paper are reproducible using a laptop with an NVIDIA GeForce RTX 3060 Laptop GPU.
@@ 2.3 Physics-Informed neural networks
and the normalization will be fulfilled if $\nu(\frac{L}{2})=1$; $\nu(-\frac{L}{2})=0$. Note that there is a trade-off here. On the one hand, the discretization error can be potentially removed using an auxiliary output. On the other hand, using it makes the network harder to train due to the additional outputs and losses. Therefore the choice to utilize an Auxiliary output or a finite difference method has to be made on a case-by-case basis.
@@ 2.5 Physical constraints and inductive biases
After defining the main differential equation loss in the previous Section, here we explain the other losses used to build a neural network capable of solving the equation. In particular, here are defined the losses linked to the physical constraints of the system. We also add a kind of loss not essential to solving the equation but which greatly improves convergence, namely the inductive biases.
To define these losses one needs to understand which are the physical constraints of the system and if any symmetries can be exploited to obtain an inductive bias. By inductive bias, we mean a set of conditions that leads the network to prioritize a set of solutions over another, for instance prioritizing solutions that follow a given symmetry. Biases can be invaluable because they limit the search space, and thus reduce the number of epochs needed to obtain a solution.
@@
where $\Omega$ is the domain of the wave function, chosen depending on the dimensionality and nature of the system. For instance, for a one-dimensional infinite potential well in $[-\frac{L}{2},\frac{L}{2}]$, $\Omega$ can be restricted to $\Omega=[-\frac{L}{2},\frac{L}{2}]$. It is important to note that imposing the condition defined in (11) is the only thing that forces the wave function to be nonzero. Otherwise, $\psi(x)=0$ for all $x$ would exactly satisfy all losses. This integral can be solved by utilizing an auxiliary output $\nu(x)$ with form (8) and by imposing $\nu(\frac{L}{2})=1$; $\nu(-\frac{L}{2})=0$. However, we must first train the network so that $\nu(x)$ has the desired form (8). Once again, this can be achieved by exploiting automatic differentiation, to compute $\gradient\nu(\vec{x})$ and thus use the following as a loss:
@@
obtaining the integral loss. We typically need to set some boundary conditions. In general, one can either use hard or soft boundary conditions. The former choice forces the network output to fulfill the boundary conditions exactly. This can be done by employing a change of variable or modifying in some way the output, such as by multiplying it by another function. The latter choice makes the network learn the BCs via an additional loss, the boundary condition loss. This loss is 00 when the BCs are fulfilled. For instance in the infinite potential well to impose soft Dirichlet Boundary conditions (10) we can add to the network’s loss $|\psi(-\frac{L}{2})|+|\psi(\frac{L}{2})|$. In this tutorial, we always use soft boundary conditions.
Further constraints can be required on a case-by-case basis, but normalization and boundary conditions are always essential. Moreover, we can also employ an additional kind of loss: the inductive biases. These are losses that are not required to obtain a solution, but can instead be added to allow the network to converge much faster. They are usually based on the symmetries of the potential: for instance, the infinite well is symmetric with respect to 00. This symmetry makes its eigenfunctions alternatively even and odd with respect to 00. A conditions that can be enforced in the PINN by the symmetry loss:
@@
where $s$ is $1$ for symmetric states and $-1$ for antysimmetric ones (the value of $s$ is switched each time a network is trained, starting from $s$=1). In general, we can utilize as inductive biases any kind of additional constraint that we know the wave function must follow. The implementation depends on the inductive bias we want to add, but in general it can be added as an additional loss. Like all neural network losses it must always be nonnegative, and must reach 00 when the bias is satisfied. It is usually effective to give such additional losses an higher weight compared to the other losses, in order to let the inductive bias better fulfill its role as a condition that limits the space of the solution. For instance in the example in Eq. (13) the total loss will greatly increase for each pair of points that does not follow the symmetry constraint. This leads the neural network to only explore the space of symmetric or anti-symmetric solutions (depending on the value of $s$).
@@ 2.6 Self-consistent computation of the eigenenergy
In this Section, we explain how to find the eigenvalue to pair with the eigenfunction obtained thanks to the techniques described in the previous Sections.
Currently, the network can only find the eigenfunction of the potential when the energy is known. However, in general when solving the Schrödinger equation that will not be the case. Therefore a way to find it in parallel with the eigenfunction is needed. There are two ways to do this: the first is to set the energy $E$ itself as a trainable parameter, but this seems to result in fairly slow training. The second is to build a second neural network to train in parallel with the one that predicts the wave function. It will take as input a constant stream of "$1"$s of the same length as the input vector, and output a guess for the energy $E$. Since the output is constant, this second network will have no hidden layers and there will be no nonlinear activation function. The network will just perform a linear transformation from $1$ to the value of the energy, based on a trainable parameter.
@@ 2.7 Finding the desired state
A crucial step is to build the losses that allow the network to distinguish the desired eigenstate from the infinite spectrum of the Hamiltonian. This is explained in this Section. By training the PINN with the current losses, it will solve the Schrödinger equation for a random state anywhere in the spectrum. Usually, however, we are looking for some amount of eigenstate starting from the ground state and going in order of increasing energy. In particular, it is especially important to build a network that finds the ground state first. This can be achieved by introducing an energy minimization condition in the loss:
@@
where $E_{PINN}$ is the predicted energy, $a$ is a hyperparameter. $E_{init}$ is either another hyperparameter when training the network to predict the ground state, or the energy of the previous state, for the excited states. The choice to increase $E_{init}$ as the quantum number increases has been made to avoid the loss exploding for higher energy states due to its exponential increase. Otherwise, the network might fail to converge at all. We also have to take into account that this energy minimization loss keeps pushing the network to lower energies even when the correct solution is reached. This moves the global minima of the loss from the eigenstate to a lower energy state, causing the loss to converge to an incorrect state. To avoid such a situation, the weight of the loss is gradually decreased as the training continues. This reduction should be such that the loss becomes approximately 00 towards the end of the training. Thanks to this in the earlier epochs this loss will steer the network towards the desired lower energy state. On the other hand in the latter epochs, the energy minimization loss will become negligible and allow the network to settle to the correct solution. The resulting neural network can now find the ground state of the wave function. We now want to obtain the excited states. To do so, a sequence of networks is built, each of which predicts one of the Hamiltonian’s eigenstates and the corresponding eigenvalue, in order of increasing energy. To do so, an additional loss is employed, the orthogonality loss. This loss forces each network’s output to be orthogonal to all the ones before it:
@@ 2.8 Weights
In the previous Sections, the losses to minimize have been defined. In the following Sections is explained how to build the actual computational framework that ensures the network is able to minimize those losses. In this Section, in particular, some pointers are given on how to tune one of the most important set of hyperparameters: the weights. Usually, when training PINNs it is better to set weights that ensure all the losses have the same scale. That is not the case for this problem: it is very easy for a network that has to predict eigenfunctions to get stuck in local minima such as $\psi(\vec{x})=0$ $\forall\vec{x}$ or an already discovered eigenvector. This means that it is actually better to set the weights in such a way that the network first enforces the physical constraints of the system, such as normalization and orthogonality. Only once these are fulfilled the network should be trained to actually solve the equation. This means setting the losses for those physical constraints higher than those of the differential equations.
In general, this means that the integral loss should have the highest weight, then the normalization, needed to avoid the constant solution, and the orthogonality, needed to avoid repeating solutions. At this point, the network should minimize the boundary conditions loss. The last loss to minimize is the differential equation loss. However, since solving the differential equation is the main objective of the network it is often worth it to gradually increase its loss’ weight. This allows the network to learn how to enforce the physical conditions in the first epochs but makes it focus more on the differential equation in the latter epochs. It is important to stress that weights are probably the most impactful hyperparameter to tune: a correct choice of weights will be the difference between quick convergence and a network getting stuck on a suboptimal solution.
@@ 2.9 Choice of training points
In order to train the PINN, the losses must be evaluated on a set of points, known as collocation points. These can be taken anywhere in the domain, and this Section contains the optimal way to choose them. When training a PINN such as this, each batch of training points should span the whole domain. Furthermore, since no dataset is being used, the choice of points is not constrained in any way: they can be taken anywhere in the domain. This can be exploited to basically work with an infinitely large dataset. To fully take advantage of this, first a batch size $N$ is decided. Then a mesh of $N$ points that spans the whole domain is built. Each of these points will be the center of a normal distribution from which one of the points in the given batch will be sampled. This results in batches that are always slightly different but always span the whole domain, thus ensuring the maximum exploration and fastest possible convergence.
@@ 2.11 Network architecture
In this Section, we give some suggestions and tricks on how to better build a feed-forward neural network to solve the Schrödinger in a physics-informed way. We focus on which metrics to use for the loss and which activation function should be employed in the network. PINNs tend to work better as relatively shallow but wide networks. For instance, in Brevi et al. (2024) the main network was made up of $7$ hidden layers of $256$ neurons each. In general, the physics-informed approach can can be applied to any kind of neural network, since its implementation is based on modifying just the loss function, independently of the network’s architecture. Here we choose to focus on feed-forward PINNs due to it being the simplest architecture, which makes it ideal for a tutorial. The number of hidden layers, of course, depends on the problem we need the PINN to address. For the examples given in this tutorial, we employed a PINN with 6 hidden layers to solve the infinite potential well and one with 10 hidden layers to find the spectrum of the particle in a ring. As for the activation function, an infinitely differentiable one such as $tanh$ is needed. Otherwise, the network would be unable to compute higher-order derivatives via automatic differentiation. Choosing an activation function of this type instead of something like $ReLU$ might, however, cause vanishing gradient problems. If this happens, it might be worth it to look into adaptive activation functions Jagtap et al. (2019) as a solution. For this tutorial, we always employ $tanh$ as the activation function. The outputs of each layer was not normalized, since empirically it was found that output normalization does not improve convergence. Lastly, it is not necessary to compute all the partial losses using the same metric. It usually makes more sense to utilize something like the Sum of Absolute Errors for the losses enforcing the physical conditions of the system, such as the Boundary Conditions loss or the Normalization loss. This is because the rest of the training makes sense only if those losses are approximately 00 in all points. And this is enforced more strictly with SAE rather than something like MSE. On the other hand, Mean Squared Error is usually the best choice for the integral and differential equation losses. Utilizing the Sum of Squared Error to compute these losses is also worth considering. This is especially true when the wave function is expected to have a complex behavior in a small section of the domain and a relatively simple one in the rest. This choice of metric will avoid underestimating the loss due to the averaging of the error for the points in the large "simple" area and the small "complex" area. As for the optimizer, Adam Kingma and Ba (2015) is usually rather effective.